OpenStudy (thefluffmuncher):

What kind of reasoning is used here? If a = b and c = d, then a – c = b – d. A. hypothesis B. conclusion C. deductive D. inductive

OpenStudy (thefluffmuncher):

im not sure but i think it's c

ganeshie8 (ganeshie8):

Yep ! you're just using the subtraction property of equality

ganeshie8 (ganeshie8):

and deducting that a-c = b-d must be true if a= b and c= d is true. so its deductive reasoning

OpenStudy (thefluffmuncher):

Which is a counterexample that disproves the following statement?

OpenStudy (thefluffmuncher):



OpenStudy (thefluffmuncher):